Simplify the expressions and solve the equations.
1. Sqrt(80b^2) -b Sqrt(125)
2. 3 Sqrt (2) + y = Sqrt(18
The first problem does'nt have an = (equal sign)
The second problem (+y) isn't square root.

Found 2 solutions by rothauserc, stanbon:
Answer by rothauserc(4718)

1. Sqrt(80b^2) -b Sqrt(125)
:
Note Sqrt(80b^2) = Sqrt(5 * 16 * b^2) and Sqrt(125) = Sqrt(5 * 25), also Sqrt(16) = + or - 4 and Sqrt(25) = + or - 5
:
***************************************
There are two solutions to this problem
:
4bSqrt(5) - 5bSqrt(5) = -bSqrt(5)
-4bSqrt(5) +5bSqrt(5) = bSqrt(5)
:
Note that (-) * (-) = +
***************************************
:
2. 3Sqrt (2) + y = Sqrt(18
y = Sqrt(18) - 3Sqrt(2)
:
**************************************
y = 3Sqrt(2) - 3Sqrt(2) = 0
y = -3Sqrt(2) - 3Sqrt(2) = -6Sqrt(2)
**************************************
:

Answer by stanbon(75887)

Simplify the expressions and solve the equations.
1. Sqrt(80b^2) -b Sqrt(125) = sqrt(5*16*b^2) - b*sqrt(5^3)
= 4b*sqrt(5) - 5b*sqrt(5)
= -b*sqrt(5)
-----

2. 3 Sqrt (2) + y = Sqrt(18
y = 3sqrt(2)-3sqrt(2)
y = 0
----------------
Cheers,
Stan H.
----------